OpenStudy (anonymous):

Solve. (x + 2)2 = 1 x = –3 and x = –1 x = 3 and x = 1 x = –1 and x = 4 x = –4 and x = –3

OpenStudy (precal):

take the square root of both sides and then subtract 2 to both sides

OpenStudy (anonymous):

(x+2)^=1 x+2=\[\pm1\] Either x+2=-1,x=-1-2=-3 Or x+2=+1,x=1-2=-1